Pde(b-c/a)yzp+(c-a/b)zxq=(a-b/c)xy
1
Expert's answer
2020-11-05T17:42:30-0500
"Solution"

"\frac {dx}{(b-c/a)yz}=\frac {dy}{(c-a/b)xz}=\frac {dz}{(a-b/c)xy} \implies""\frac {xdx+ydy+zdz}{xyz(b-c/a+c-a/b+a-b/c)}=\frac {dx}{(b-c/a)yz}"

"\frac {x^2+y^2+z^2}{b-c/a+c-a/b+a-b/c}=\frac {x^2}{b-c/a}+C_1"


"\frac {xdx+ydy+zdz}{xyz(b-c/a+c-a/b+a-b/c)}=\frac {dy}{(c-a/b)xz}"

"\frac {x^2+y^2+z^2}{b-c/a+c-a/b+a-b/c}=\frac {y^2}{c-a/b}+C_2"


"\frac {xdx+ydy+zdz}{xyz(b-c/a+c-a/b+a-b/c)}=\frac {dz}{(a-b/c)xy}"

"\frac {x^2+y^2+z^2}{b-c/a+c-a/b+a-b/c}=\frac {z^2}{a-b/c}+C_3"